Write the equation, in standard form, that passes through (7,2) and (-1,6).
Answer by checkley75(3666)

first we need to find the slope (m)
(y2-y1)/(x2-x1)
(6-2)/(-1-7)
4/-8
-1/2 is the slope for this line.
now substitute one set of x,z values & solve for the y intyercept(b)
2=-1/2*7+b
2=-7/2+b
b=2+7/2
b=2+3.5
b=5.5
thus the line equation is
y=-1/2x+5.5